# Rigidity of Gradient Shrinking Ricci Solitons with Bach-like Flatness and Related Variational Formulas

### v2

@@ Abstract
The classical Bach tensor in four dimensions can be expressed as a linear combination of two independent, symmetric, divergence-free, quadratic in curvature tensors U and V. Several classification results for gradient-shrinking Ricci solitons have been obtained under the assumption that the Bach tensor vanishes. We define a Bach-like tensor to be any other linear combination of U and V. We prove that, except along the Bach line, the vanishing of a Bach-like tensor forces a four-dimensional complete gradient-shrinking Ricci soliton to be either Einstein or isometric to the Gaussian soliton. Finally, we show that Bach-like tensors arise as Euler–Lagrange equations of a two-parameter family of quadratic curvature functionals and compute the corresponding first and second variation formulas.
@@ 1 Introduction and Main Results
for $\rho$ a constant. The soliton is said to be shrinking if $\rho>0$, and it is common to scale $f$ so that $\rho=\frac{1}{2}$ for convenience. Such solitons arise as self-similar solutions of Hamilton’s Ricci flow, as models for type I singularities. Unless otherwise stated, our work will be in dimension $n=4$.
The classical Bach tensor was introduced by R. Bach [bach:1921] to study conformal relativity. It is defined in dimension 4 by:
@@
where $\mathrm{W}_{ijkl}$ denotes the Weyl curvature and $R_{ij}$ the Ricci curvature. It is well-known that $\mathrm{B}_{ij}$ is conformally invariant, symmetric, trace-free, divergence-free, and quadratic in the Riemann curvature tensor. As seen in [bergman:2004] there are exactly two independent symmetric 22-tensors that are divergence-free and quadratic in curvature in four dimensions. These are the following:
@@
One can express the Bach tensor as a combination of $U$ and $V$. In dimension 4 we have:
@@
for constants $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. If $\alpha/\beta=3$ then the Bach-like tensor is simply a rescaled Bach tensor. H.-D. Cao and Q. Chen showed in [caochen:2013] that if the Bach tensor vanishes on a complete gradient-shrinking Ricci soliton, then it must be rigid. Once we move beyond the Bach tensor itself, we can show that Bach-like flatness still implies rigidity, but without the index-heavy computations typical of the Bach-flat case. On the one hand, our results here are somewhat stronger in that only two classifications occur; on the other, this suggests that the Bach tensor behaves more like a genuine analytic obstruction, requiring more delicate analysis that ties directly into the geometry of the conformal Cotton tensor $\mathrm{D}_{ijk}$, which will be discussed later.
Theorem 1.1 (Vanishing Basis-Tensors).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then
  1. (i)
    $\mathrm{V}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton;
  2. (ii)
    $\mathrm{U}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton.
@@
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Assume $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$ such that $\beta\neq\alpha/3$. Then if
$\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0,$
$M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton. Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathfrak{B}_{ij}=0$. Hence, within the class of complete gradient-shrinking Ricci solitons, $\mathfrak{B}_{ij}=0$ is equivalent to $M^{4}$ being Einstein or isometric to the Gaussian soliton.
We will approach the proof of Theorem 1.2 by first considering the cone
@@
and primarily using weighted-integral identities from the $\mathrm{V}_{ij}$ tensor. For the region outside the cone we will take a different approach using more general weighted-integral identities. This result covers the entire $(\alpha,\beta)$ plane aside from the Bach-line and the origin.
@@ 2 Integral Identities
To proceed, we rely on a number of integral identities implied by $\Delta R=0$. The connection between harmonic scalar curvature and Bach-like tensors is rooted in the traces of U and V.
@@
Let $U_{ij}$ and $V_{ij}$ be given by (3) and (4) respectively. Then
$\mathrm{tr}(\mathrm{U}_{ij})$ $=-\Delta R$
$\mathrm{tr}(\mathrm{V}_{ij})$ $=3\Delta R.$
So if $(\alpha,\beta)\in\mathcal{C}$, then $\mathfrak{B}_{ij}=0$ implies $\Delta R=0.$ We now derive the integral identities for both the unweighted and weighted volume elements.
@@
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and assume that $\Delta\mathrm{R}=0$. Let $r$ be a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. Then:
@@
$0=\int_{\Omega_{r}}\Delta R\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla R,\nabla f\rangle\,\mathrm{d}S$
@@
As $\Delta R=0$ by assumption, and using (1) we see that both integrals must vanish, showing (2). To show (3) we integrate $R\Delta R$:
@@
And by using the identity $\mathrm{R}=f-|\nabla f|^{2}$ from normalizing $f$ as in [hamilton:1988], we get (3). Now we have:
@@
showing (4). To show (5) we integrate $R\Delta R$ with the weighted volume $\mathrm{e}^{-f}\,\mathrm{d}V$:
@@
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-R\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
@@ 3 The Vanishing V-tensor
The tensor $\mathrm{V}_{ij}$ has fewer and less complicated terms than $\mathrm{U}_{ij}$. We will obtain integral identities by assuming $\mathrm{V}_{ij}=0$, and use these to prove Theorem 1.1. We first consider the Hessian term:
@@
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and assume that $\Delta R=0$. Then if $r$ is a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$, we have:
@@
$\frac{1}{4}|\nabla\mathrm{R}|^{2}=|\mathrm{Rc}(\nabla f)|^{2}\leq|\mathrm{Rc}|^{2}|\nabla f|^{2}$
By H.-D. Cao and D. Zhou’s growth estimates for the potential function [caozhou:2010], we have:
$|\nabla f|^{2}(x)\leq\frac{1}{4}(d(x)+c_{2})^{2}$
@@
Clearly the first integral over the compact set $K$ is finite. Now we have for the second integral:
@@
By O. Munteanu and N. Sesum’s results in [munteanusesum:2013], we have:
@@
$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V<\infty$
@@
And using Lemma 3.1, we compute the integral of $V(\nabla f,\nabla f)$ with respect to the weighted volume element $\mathrm{e}^{-f}\,\mathrm{d}V$, and we note that the second term involving $\Delta R$ again vanishes:
@@
Assuming $\mathrm{V}_{ij}=0$, for all $p\in M$ we must have either $\mathrm{R}=0$ or $|\nabla f|=0$. By the work of S. Pigola, M. Rimoldi, and A. Setti [pigolarimoldi:2011], either $R>0$ everywhere, hence $\nabla f=0$ and the manifold is Einstein, or if $R=0$ anywhere, then the soliton must be isometric to the Gaussian soliton on $\mathbb{R}^{4}$.
@@
$\mathrm{V}_{ij}$ $=\frac{1}{4}\mathrm{R}^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}$
$=-\mathrm{R}(\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}\,g_{ij})$
$=0.$
@@ 5 Weighted Bach Integral
In [caochen:2013] H.-D. Cao and Q. Chen consider how the Bach tensor links to the Cotton tensor of a conformal metric, called $\mathrm{D}_{ijk}$. The norm of $\mathrm{D}_{ijk}$ contains rich geometric information, and they were able to use it to prove their classification results. We will link Bach-like tensors to $\mathrm{D}_{ijk}$ as well. Recall that the Cotton tensor is given by:
@@
where $\mathrm{W}_{ijkl}$ is the Weyl curvature. The Cotton tensor is skew-symmetric in the first two indices and trace-free in any two indices. That is,
@@
$D_{ijk}$ $=\frac{1}{n-2}\big(\mathrm{R}_{jk}\nabla_{i}f-\mathrm{R}_{ik}\nabla_{j}f\big)+\frac{1}{2(n-1)(n-2)}\big(g_{jk}\,\nabla_{i}\mathrm{R}-g_{ik}\,\nabla_{j}\mathrm{R}\big)$
@@
H.-D. Cao and Q. Chen [caochen:2013] showed how the Bach tensor can be expressed in terms of $\mathrm{C}_{ijk}$ and $\mathrm{D}_{ijk}$:
@@
We can now link together the integral of the Bach tensor contracted with the gradient of the potential function to the norm of $\mathrm{D}_{ijk}$.
@@
Let $r$ be a regular value of $f$ and $\Omega_{r}=\Omega_{r}=\{f\leq r\}$. We use the antisymmetric properties of the Cotton and $\mathrm{D}_{ijk}$ tensor:
@@
First observe that for any $(\alpha,\beta)\in\mathcal{C}$, and a Bach-like tensor $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}$ we can write:
@@
Since $(\alpha,\beta)\in\mathcal{C}$, we have two cases. If $\alpha>0$ and $\beta>\alpha/3$, then we must have $|\mathrm{D}_{ijk}|^{2}(p)=0$ and $\mathrm{R}^{2}(p)|\nabla f(p)|^{2}=0$ for all $p\in M$. Then we can conclude that $M$ is either Einstein or isometric to the Gaussian soliton as in the proof of Theorem 1.1. The argument for when $\alpha<0$ and $\beta<\alpha/3$ is analogous.

### v3

# arXiv:2511.05774v3[math.DG] 02 Mar 2026

@@ Abstract
The classical Bach tensor in four dimensions can be expressed as a linear combination of two independent, symmetric, divergence-free, quadratic-in-curvature tensors $\mathrm{U}$ and $\mathrm{V}$. Several classification results for gradient-shrinking Ricci solitons have been obtained under the assumption that the Bach tensor vanishes. We define a Bach-like tensor to be any other linear combination of $\mathrm{U}$ and $\mathrm{V}$. We prove that within a certain cone of parameters $\mathcal{C}$, the vanishing of a Bach-like tensor forces a four-dimensional complete gradient-shrinking Ricci soliton to be either Einstein or isometric to the Gaussian soliton, extending the results of Cao–Chen [4]. The special case where $\mathrm{U}=0$ forces $f_{\min}\in\{0,1,2\}$, with rigidity holding when $f_{\min}=0,2$. The remaining case $f_{\min}=1$ is the central open problem, with $\mathbb{R}^{2}\times\mathbb{S}^{2}(\sqrt{2})$ as the conjectured exceptional geometry. Finally, we show that Bach-like tensors arise as Euler–Lagrange equations of a two-parameter family of quadratic curvature functionals and compute the corresponding first and second variation formulas.
@@ 1 Introduction and Main Results
for $\rho$ a constant. The soliton is said to be shrinking if $\rho>0$, and it is common to scale $f$ so that $\rho=\frac{1}{2}$ for convenience. We normalize $f$ throughout so that
$\mathrm{R}+|\nabla f|^{2}=f,$ (2)
following Hamilton [6]. Such solitons arise as self-similar solutions of Hamilton’s Ricci flow, as models for type I singularities. Unless otherwise stated, our work will be in dimension $n=4$.
The classical Bach tensor was introduced by R. Bach [2] to study conformal relativity. It is defined in dimension 4 by:
@@
where $\mathrm{W}_{ijkl}$ denotes the Weyl curvature and $\mathrm{R}_{ij}$ the Ricci curvature. It is well-known that $\mathrm{B}_{ij}$ is conformally invariant, symmetric, trace-free, divergence-free, and quadratic in the Riemann curvature tensor. As seen in [1] there are exactly two independent symmetric $2$-tensors that are divergence-free and quadratic in curvature in four dimensions:
@@
One can express the Bach tensor as a combination of $\mathrm{U}$ and $\mathrm{V}$. In dimension 4:
@@
for constants $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. If $\alpha/\beta=3$ then the Bach-like tensor is simply a rescaled Bach tensor. H.-D. Cao and Q. Chen showed in [4] that if the Bach tensor vanishes on a complete gradient-shrinking Ricci soliton, then it must be rigid. Once we move beyond the Bach tensor itself, we can show that Bach-like flatness still implies rigidity within a certain parameter cone, but without the index-heavy computations typical of the Bach-flat case. The principal results of this paper are as follows.
Theorem 1.1 (Vanishing V-tensor).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Then $\mathrm{V}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton. In particular, $f_{\min}\in\{0,2\}$ correspondingly.
Theorem 1.2 (Vanishing U-tensor).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). If $\mathrm{U}_{ij}\equiv 0$ then $f_{\min}\in\{0,1,2\}$. If $f_{\min}=0$ then $M^{4}$ is isometric to the Gaussian shrinking soliton, and if $f_{\min}=2$ then $M^{4}$ is Einstein. Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathrm{U}_{ij}\equiv 0$.
The case $f_{\min}=1$ in Theorem 1.2 is not resolved by the methods of this paper. At the minimum point $p$ of $f$, the tensor $\mathrm{U}$ vanishes independently of any global assumption when $\mathrm{R}(p)=1$, so the local analysis yields no information. By contrast, $\mathrm{V}$ cannot vanish at $p$ when $f_{\min}=1$. We state the following as a conjecture.
Conjecture 1.3.
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). If $\mathrm{U}\equiv 0$ and $f_{\min}=1$, then $M^{4}$ is isometric to $\mathbb{R}^{2}\times S^{2}(\sqrt{2})$.
The candidate space $\mathbb{R}^{2}\times S^{2}(\sqrt{2})$ is verified to satisfy the hypotheses. For the shrinking soliton normalization $\rho=\tfrac{1}{2}$, the soliton equation in the $S^{2}$ directions forces radius $r=\sqrt{2}$, giving $\mathrm{R}=1$ and $f_{\min}=1$ via (2). Indexing by $0,1$ in the $\mathbb{R}^{2}$-directions and $a,b$ on $S^{2}(\sqrt{2})$, a direct computation gives $\mathrm{U}_{00}=\mathrm{U}_{0a}=\mathrm{U}_{ab}=0$, so $\mathrm{U}\equiv 0$. Moreover, $\mathrm{V}_{00}=-\tfrac{1}{4}$ and $\mathrm{V}_{ab}=\tfrac{1}{4}g_{ab}$, so $\mathrm{V}\not\equiv 0$, confirming the space is neither Einstein nor Gaussian. This computation was suggested to us by H.-D. Cao.
@@
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Assume $(\alpha,\beta)\in\mathcal{C}$ where
@@
and assume $3\beta\neq\alpha$. If $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$, then $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton. Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathfrak{B}_{ij}=0$. If $3\beta=\alpha$, the rigidity results of Cao–Chen [4] apply.
We approach the proof of Theorem 1.4 by primarily using weighted integral identities from the $\mathrm{V}_{ij}$ tensor inside the cone $\mathcal{C}$. This result covers the entire $(\alpha,\beta)$ plane aside from the Bach line, the $\beta=0$ axis, and the origin; a complete treatment of the remaining cases is contingent on Conjecture 1.3 and is deferred to future work.
@@ 2 Integral Identities
To proceed, we rely on a number of integral identities implied by $\Delta\mathrm{R}=0$. The connection between harmonic scalar curvature and Bach-like tensors is rooted in the traces of $\mathrm{U}$ and $\mathrm{V}$.
@@
Let $\mathrm{U}_{ij}$ and $\mathrm{V}_{ij}$ be given by (4) and (5) respectively. Then
$\mathrm{tr}(\mathrm{U}_{ij})$ $=-\Delta\mathrm{R},$
$\mathrm{tr}(\mathrm{V}_{ij})$ $=3\Delta\mathrm{R}.$
So if $(\alpha,\beta)\in\mathcal{C}$, then $\mathfrak{B}_{ij}=0$ implies $\Delta\mathrm{R}=0.$ We now derive the integral identities for both the unweighted and weighted volume elements.
@@
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2), and assume that $\Delta\mathrm{R}=0$. Let $r$ be a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. Then:
@@
$0=\int_{\Omega_{r}}\Delta\mathrm{R}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S$
@@
As $\Delta\mathrm{R}=0$ by assumption, and using (1) we see that both integrals must vanish, showing (2). To show (3) we integrate $\mathrm{R}\Delta\mathrm{R}$:
@@
And by using the normalization (2), we get (3). Now we have:
@@
showing (4). To show (5) we integrate $\mathrm{R}\Delta\mathrm{R}$ with the weighted volume $\mathrm{e}^{-f}\,\mathrm{d}V$:
@@
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
@@ 3 The Vanishing V-tensor
The tensor $\mathrm{V}_{ij}$ has fewer and less complicated terms than $\mathrm{U}_{ij}$. We will obtain integral identities by assuming $\mathrm{V}_{ij}=0$, and use these to prove Theorem 1.1. We first consider the Hessian term.
@@
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2), and assume that $\Delta\mathrm{R}=0$. Then if $r$ is a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$, we have:
@@
$\frac{1}{4}|\nabla\mathrm{R}|^{2}=|\mathrm{Rc}(\nabla f)|^{2}\leq|\mathrm{Rc}|^{2}|\nabla f|^{2}.$
By H.-D. Cao and D. Zhou’s growth estimates for the potential function [5], we have:
$|\nabla f|^{2}(x)\leq\frac{1}{4}(d(x)+c_{2})^{2},$
@@
The first integral over the compact set $K$ is finite. We show that the second integral is also finite:
@@
By O. Munteanu and N. Sesum’s results in [7], we have:
@@
$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V<\infty.$
@@
And using Lemma 3.1, we compute the integral of $\mathrm{V}(\nabla f,\nabla f)$ with respect to the weighted volume element $\mathrm{e}^{-f}\,\mathrm{d}V$, noting that the term involving $\Delta\mathrm{R}$ vanishes:
@@
Assuming $\mathrm{V}_{ij}=0$, we must have either $\mathrm{R}=0$ or $|\nabla f|=0$ pointwise. By the work of S. Pigola, M. Rimoldi, and A. Setti [8], either $\mathrm{R}>0$ everywhere, hence $\nabla f=0$ and the manifold is Einstein, or if $\mathrm{R}=0$ anywhere, then the soliton must be isometric to the Gaussian soliton on $\mathbb{R}^{4}$. The values $f_{\min}\in\{0,2\}$ follow from (2): if Einstein then $\nabla f\equiv 0$ and $\mathrm{Rc}=\tfrac{1}{2}g$, giving $\mathrm{R}=2$ and $f_{\min}=2$; if Gaussian then $\mathrm{R}=0$ and $f_{\min}=0$.
@@
$\mathrm{V}_{ij}$ $=\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}^{2}g_{ij}=-\mathrm{R}\Big(\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}\,g_{ij}\Big)=0.$
@@ 5 Weighted Bach Integral
In [4] H.-D. Cao and Q. Chen consider how the Bach tensor links to the Cotton tensor of a conformal metric, called $\mathrm{D}_{ijk}$. The norm of $\mathrm{D}_{ijk}$ contains rich geometric information, and they were able to use it to prove their classification results. We will link Bach-like tensors to $\mathrm{D}_{ijk}$ as well. Recall that the Cotton tensor is given by:
@@
where $\mathrm{W}_{ijkl}$ is the Weyl curvature. The Cotton tensor is skew-symmetric in the first two indices and trace-free in any two indices:
@@
$\mathrm{D}_{ijk}$ $=\frac{1}{n-2}\big(\mathrm{R}_{jk}\nabla_{i}f-\mathrm{R}_{ik}\nabla_{j}f\big)+\frac{1}{2(n-1)(n-2)}\big(g_{jk}\,\nabla_{i}\mathrm{R}-g_{ik}\,\nabla_{j}\mathrm{R}\big)$
@@
H.-D. Cao and Q. Chen [4] showed how the Bach tensor can be expressed in terms of $\mathrm{C}_{ijk}$ and $\mathrm{D}_{ijk}$:
@@
Let $r$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. We use the antisymmetric properties of the Cotton and $\mathrm{D}_{ijk}$ tensor:
@@
First observe that for any $(\alpha,\beta)\in\mathcal{C}$, we can write:
@@
Since $(\alpha,\beta)\in\mathcal{C}$, if $\alpha>0$ and $\beta>\alpha/3$ then both integrals are nonnegative and must each vanish, so $|\mathrm{D}_{ijk}|^{2}\equiv 0$ and $\mathrm{R}^{2}|\nabla f|^{2}\equiv 0$. We conclude that $M$ is either Einstein or isometric to the Gaussian soliton as in the proof of Theorem 1.1. The argument for $\alpha<0$ and $\beta<\alpha/3$ is analogous.